A 0.40 M solution of C3H5O3- has a pH of 8.728. What is the Kb of this ion?


1
Expert's answer
2021-04-25T07:40:22-0400

Solution:

For any aqueous solution at 25C:

pH + pOH = 14

pOH = 14 - pH = 14 - 8.728 = 5.272

We can convert between pOH and [OH] using the following equation:

pOH = −log[OH]

[OH] = 10−pOH = 10−5.272 = 5.346×10−6 M

[OH] = 5.346×10−6 M


Consider the equilibrium,

C3H5O3 + H2O ⇌ HC3H5O3 + OH

where,




From this, we can construct an ICE(Initial, Change, Equilibrium) Table:



Hence,

Kb = [HC3H5O3] × [OH] / [C3H5O3]

Kb = (5.346×10−6 M) × (5.346×10−6 M) / (0.399995 M) = 7.145×10−11 M

Kb = 7.145×10−11


Answer: Kb = 7.145×10−11